OpenStudy (anonymous):

Simple geometry/algebra question! I just want to make sure I'm doing this right… What number is a solution to x/12=x-3/8 a. 2 b. 3 c. 6 d. 9 (what I got) Thanks! x

OpenStudy (anonymous):

Yup you got it right, just double-checked it :)

OpenStudy (anonymous):

@ajm r u there?

OpenStudy (anonymous):

Also to confirm the answer plug-in 9 for x 9/12=6/8= 3/4 So correct :)

OpenStudy (anonymous):

Thanks so much!